# Probing entanglement dynamics and topological transitions on noisy intermediate-scale quantum computers

### v2

@@ I Introduction
In this work, we simulate quench dynamics of the Su-Schrieffer-Heeger (SSH) chain [20], a paradigmatic model for topological insulators, on IBM-Q devices. Considerable efforts have been made over the past years to elucidate topological dynamical phenomena in quantum quenches between different topological phases in the SSH model, predominantly focusing on the Bloch state in momentum space [21, 22, 23, 24, 25]. Here we focus on observables in real space, for up to 12 qubits. We quantify the entanglement using the second-order Rényi entanglement entropy and probe topological properties using the twist operator appearing in the Lieb-Schultz-Mattis theorem [26, 27]. We consider a special case where the dynamics of entanglement and topology exhibits persistent and periodic oscillations. This simulation can be executed without Trotter errors, enabling us to evaluate the ability of NISQ devices to simulate quantum many-body dynamics through comparisons with exact solutions.
The paper is organized as follows: In Sec. II, we define the model and the quench protocols considered in our simulations; we summarize some known properties of the quench dynamics. In Sec. III, we describe our simulation setups on the IBM-Q devices and present our simulation results, focusing on the Rényi entanglement entropy and twist order parameters. We conclude in Sec. IV with a summary and discussion.
@@ II The Model and quench protocols
where $m$ indexes $N$ unit cells that are pairs of two sites, one on sublattice $A$ and one on $B$. The operators $a^{\dagger}_{m}$($b^{\dagger}_{m}$) creates a spinless fermion at the $A$($B$)-sublattice site in the $m$th unit cell, and $\mathcal{N}$ is the the number of fermions. The constants $2J$ and $2J^{\prime}$ are regarded as the intracell and intercell hopping amplitudes, respectively. The boundary term is affected by the parity of the number of fermions [26]; depending on evenness or oddness of $\mathcal{N}$, the boundary conditions of the fermionic chain are antiperiodic or periodic.
@@
which is equivalent to the ground state of the trivial flat-band Hamiltonian with $(J,J^{\prime})=(1,0)$. For a quench from $\left|{\psi_{\text{N}}}\right\rangle$, the post-quench state will reach a winding number $\nu=1$ at times $t_{\text{N}}=(n-\frac{1}{2})\frac{\pi}{4}$ with $n\in\mathbb{Z}^{+}$; for a quench from $\left|{\psi_{\text{S}}}\right\rangle$, the post-quench state will have a winding number $\nu=2$ at $t_{\text{S}}=(n-\frac{1}{2})\frac{\pi}{2}$ [21, 22]. Below we denote by $t_{\text{N}}^{*}$ and $t_{\text{S}}^{*}$ those time points at which a finite winding number occurs. In real space, the periodic generation of the winding number manifests itself in the dynamics of intercell entanglement formation [33]. For the initial $\left|{\psi_{\text{N}}}\right\rangle$ state, the post-quench state consists of a set of nearest neighbor intercell entangled states at $t=t_{\text{N}}^{*}$; For the initial $\left|{\psi_{\text{S}}}\right\rangle$ state, the post-quench state develops entangled states connecting $A$ and $B$ sublattices in two next-nearest-neighbor cells (i.e. extended to three lattice spacings) at times $t=t_{\text{S}}^{*}$.
@@ II.1 Entanglement entropy
We note that $S^{0\to 2}(t)=2S^{0\to 1}(t/2)$ [see Fig. 1(a)]. For a short closed chain of length $L=4$ with only two unit cells, the dynamical Rényi EE is reduced to $S^{0\to 1}(t)$.
Figure 1: Exact data for quenches from a Néel state $\left|{\psi_{\text{N}}}\right\rangle$ (solid lines) and from a fully dimerized state $\left|{\psi_{\text{S}}}\right\rangle$ (dashed lines), driven by the flat-band Hamiltonian with $(J=0,J^{\prime}=1)$ and periodic boundary conditions. The panels from top to bottom show the evolution of (a) the Rényi entanglement entropy with a symmetric bipartition, (b) the twist order parameter for $L=16$ and, (c) the Berry phase for $L=16$.
@@ II.2 Twist operator
for a spin chain of length $L$, is a useful tool to characterize various topology properties. This unitary operator was first introduced in the Lieb-Schultz-Mattis (LSM) theorem [26], which states that the amplitude, dubbed as the twist order parameter, given by
@@
in terms of the ground state of the XXZ spin chain can be seen as an indicator of a gapped or gapless excitation spectrum, depending on whether $z^{(q)}_{L}$ is nonvanishing or zero in the thermodynamic limit $L\to\infty$. Furthermore, the twist order parameter was generalized to higher spin cases and is capable of detecting the patterns of valence-bond-solid-like dimerization in the state $\left|{\psi}\right\rangle$ [27], even for random and inhomogeneous systems [35]. The real part of $z_{L}$ is negative/positive for a dimerized state with strong links connecting adjacent sites $(j,j+1)$ with even/odd $j$ (i.e. intercell/intracell links).
Here we apply the twist operator to the dynamical states in our quench protocols. As shown in Fig. 1(b) for $L=16$, the twist order parameter oscillates in the region $\text{Re}[z_{L}]\leq 0$ for a quench from the Néel state $\left|{\psi_{\text{N}}}\right\rangle$, while it oscillates in the region $\text{Re}[z_{L}]>0$ for a quench from the initial state $\left|{\psi_{\text{S}}}\right\rangle$. The oscillation of $\text{Re}[z_{L}]$ for the case of $\left|{\psi_{\text{N}}}\right\rangle$ indicates that the strongest dimerization occurs at $t^{*}_{\text{N}}$ when the maximum entanglement forming between nearest neighboring cells; for the case of $\left|{\psi_{\text{S}}}\right\rangle$, the system starts from a fully dimerized state with intracell links and returns to this strongest dimerization after a period of $\pi/2$.
For the free-fermion model we consider here, the twist operator can be used to define localization length and electric polarization [36, 37]. Following Refs. [36, 38], we consider the unitary operator in terms of the particle number operator, $n_{j}=S^{z}_{j}+1/2$, at each site,
$U_{n}=\exp\biggl{(}i\frac{2\pi}{L}\sum_{j=1}^{L}jn_{j}\biggr{)}\,.$ (16)
Similar to the twist order parameter $z_{L}$, the ground-state expectation value
@@
is an order parameter that distinguishes an insulating phase and a conducting phase at rational filling $n=p/q$ (where $p/q$ is an irreducible fraction) [39], depending on whether $z^{(q)}_{N}\neq 0$ or $z^{(q)}_{N}=0$ in the limit of $L\to\infty$ [37]. The argument of $z^{(q)}_{N}$ defines the center of mass of the particles, which is related to the Berry phase $\gamma_{L}$ via [36, 39]
$\gamma_{L}=\text{Im}\ln[z^{(q)}_{N}]\,.$ (18)
We apply $U^{q}_{n}$ with $q=2$ for our dynamical states at half filling ($n=1/2$). As shown in Fig. 1(c) The Berry phase for $\left|{\psi_{\text{N}}(t)}\right\rangle$ jumps between $-\pi$ and $\pi$ at $t=t_{\text{N}}^{*}$, indicating a quantized dynamical Chern number $C=\pm 1$; this result agrees with previous studies based on the single-particle Hamiltonian in momentum space [25]. On the other hand, the Berry phase for $\left|{\psi_{\text{S}}(t)}\right\rangle$ remains zero for all $t$, reflecting that the dynamical state has inversion symmetry.
@@ III.1 Time evolution
The default state on the IBM devices is a tensor product of $\left|{0}\right\rangle$ and the computational basis is also composed by the tensor product of the eigenstates of $\sigma^{z}$. We set up the initial states $\left|{\psi_{\text{N}}}\right\rangle$ and $\left|{\psi_{\text{S}}}\right\rangle$ for our quench protocol using the Pauli gate $\bm{X}$, the Hadamard gate $\bm{H}$ and the controlled-not (CNOT) gate $\bm{C_{X}}$, as shown in Fig. 2, where a two-spin singlet state is realized by a set of quantum gates $(\bm{X}\otimes\bm{X})(\bm{H}\otimes\bm{I})(\bm{C_{X}})$.
@@
For the post-quench Hamiltonian, the fully dimerized XX chain with $(J,J^{\prime})=(0,1)$, that we consider here, the quantum circuit for the time evolution operator $U(t)=\exp(-iHt)$ with an arbitrary time duration $t$ can be constructed without the use of the Trotter decomposition. The time evolution operator can be split into
@@
where the commutator identity $[\sigma_{j}^{x}\sigma_{j^{\prime}}^{x},\,\sigma_{j}^{y}\sigma_{j^{\prime}}^{y}]=0$ is applied. The quantum circuits required for the time evolution are then the gates to implement the operators $U^{xx}(t)=e^{-it\bm{X}\bm{X}}$ and $U^{yy}(t)=e^{-it\bm{Y}\bm{Y}}$, which can be transformed from the time evolution operator of an Ising spin pair $U^{zz}(t)=e^{-it\bm{Z}\bm{Z}}$. The circuit for the operator $U^{zz}(t)$ is composed of a rotation gate $\bm{R_{z}}(2t)$ and two $\bm{C_{X}}$ gates, as shown in Fig. 3 (a). We can convert the operator $U^{zz}(t)$ into $U^{xx}(t)$ using Hadamard gates via $\bm{H}\bm{Z}\bm{H}=\bm{X}$. Similarly, we implement $U^{yy}(t)$ by wrapping the time-evolution operator $U^{zz}(t)$ with Hadamard gates and $\bm{S}$-gates on each qubit, where $\bm{S}$ is a phase shift gate that rotates a $\pi/2$ radian about the $z$-axis, inducing $\bm{S}\bm{X}\bm{S}=\bm{Y}$. The circuits for $U^{xx}(t)$ and $U^{yy}(t)$ are shown in Fig. 3 (b) and Fig. 3 (c), respectively.
@@ III.2 The Rényi entanglement entropy
The key ingredient of the second-order Rényi entropy is the purity $\text{Tr}[\rho^{2}]$ of a given (reduced) density matrix. A recently developed measurement protocol employing statistical correlations of locally randomized measurements [40, 41, 42] enables us to measure the purity with an adequate accuracy on current noisy quantum devices via classical postprocessing. This measurement procedure for our simulation of the Rényi entanglement entropy is as follows: (i) Generate a post-quench state, $\rho(t)$, of $L$ qubits at time $t$; (ii) Draw $N_{U}$ sets random unitaries of the form $U=u_{1}\otimes\cdots\otimes u_{L}$, where the $u_{i}$ are single-qubit Haar random unitaries sampled independently from the circular ensemble [41, 43]; (iii) Apply each set of $U$ to the state $\rho(t)$ and measure the the probabilities $P(\mathbf{s}_{I})$ of obtaining computational basis states $\left|{\mathbf{s}_{I}}\right\rangle=\left|{s_{1},\cdots,s_{N_{I}}}\right\rangle$ for the considered subsystem $I$ of $N_{I}$ qubits. The purity $\text{Tr}[\rho_{I}^{2}]$ for the reduced density matrix is then estimated using
@@
Figure 5: Results for the half-chain entanglement entropy for a chain with $L=4$ and PBC after quenches from the initial states $\left|{\psi_{\text{N}}}\right\rangle$ and $\left|{\psi_{\text{S}}}\right\rangle$, simulated on the IBM device ibm_mumbai.
@@
Figure 6: Results for the half-chain entanglement entropy for a chain with $L=8$ and PBC after quenches from the initial states $\left|{\psi_{\text{N}}}\right\rangle$ (a) and $\left|{\psi_{\text{S}}}\right\rangle$ (b), simulated on the IBM device ibm_montreal. The results for $\left|{\psi_{\text{S}}}\right\rangle$ after applying error mitigation using the global depolarizing ansatz show an excellent match with the exact solution.
@@
For longer chains, there is however a serious decline in the accuracy of the simulation results, as shown in Fig. 6 for $L=8$. For the initial state $\left|{\psi_{\text{N}}}\right\rangle$, there is a entropy shift when compared to the exact solution. This offset, possibly some inherent error in the device, can be partially eliminated by aligning the entropy at $t=0$ with the exact value $S(t=0)=0$. We note that the shift mitigation approach has also applied in previous studies [45]. To eliminate the deviations beyond the constant shift, we apply a simple error mitigation strategy using the global depolarizing noise ansatz, along the lines of Ref. [15]. We assume that the entire quantum state of $L$ qubits under the noise is
@@
where $I$ denotes a subsystem of $N_{I}$ qubits. The error rate $p_{\text{tot}}$ can be extracted from $\text{Tr}[\rho^{2}]$ when the full density matrix for a pure state is considered since $\text{Tr}[\rho^{2}_{\text{exact}}]=1$. We have determined $p_{\text{tot}}$ for each time step to carry out the error mitigation for the time-dependent data. In addition, the mitigated data in Fig. 6 are adjusted to align the minimum to zero to correct negative entropies in the minima. The results for $\left|{\psi_{\text{S}}}\right\rangle$ after applying error mitigation clearly show an excellent match with the exact data, while the mitigated results for $\left|{\psi_{\text{N}}}\right\rangle$ with global depolarizing ansatz are merely passable.
Figure 8: A schematic plot of the heavy-hexagon lattice, widely adopted in the IBM quantum processors. Each circle represents a physical qubit. The blue circles form a ring layout and the red circles form a line layout that are selected in our simulations for a 8-qubit chain and a 12-qubit chain, respectively.
@@
The deviation from the exact solutions can have more than one source. For example, the accuracy of the estimated purity by randomized measurements depends crucially on the the number of random unitaries and the number of measurements per random unitary [41], as implied in Eq. (21). However, the primary source of error for our case may stem from the PBC imposed on the model, which requires connectivity between the first and last qubits via ancilla qubits and additional gates, thereby leading to higher error rates. For example, the circuit depth for the simulations of the chains with PBC is 40 (45) for the initial state $\left|{\psi_{\text{N}}}\right\rangle$ ($\left|{\psi_{\text{S}}}\right\rangle$), while the depth is reduced to 15 (19) when open chains with OBC are considered. We can indeed reproduce the results for $L=8$ with higher accuracy by using a segment of bulk qubits (from the 3rd to the 6th qubit) in an open chain as the subsystem considered for the EE. The results for the open chain displayed in Fig. 7 largely agree quantitatively with the exact data, even without applying error mitigation. After applying the global depolarizing error mitigation, the overall shape of the EE as a function of time, in particular around the entropy valleys, is corrected. However, we need to apply a constant shift to align the entropy at $t=0$ (the minimum entropy) to zero.

### v3

@@ I Introduction
In this work, we simulate quench dynamics of the Su-Schrieffer-Heeger (SSH) chain [20] on IBM-Q devices. The SSH model serves as a fundamental system for describing a topological insulator that exhibits transitions between topologically distinct phases in its quench dynamics [21, 22, 23, 24]. Previous simulations of the SSH model and related models on NISQ devices have predominantly focused on Bloch states in momentum space [25, 26, 27] or edge states [28, 29]. Here we focus on topological properties of the entire system in real space, with up to 12 qubits. We quantify the entanglement using the second-order Rényi entanglement entropy and probe topological properties using the twist operator appearing in the Lieb-Schultz-Mattis theorem [30, 31]. We consider a special case where the dynamics of entanglement and topology for an arbitrary system size exhibits persistent and periodic oscillations. We obtain analytical closed-form solutions for the dynamical entanglement entropy starting from two different initial states. The simulations can be executed without Trotter errors and without using variational approaches, enabling us to evaluate the ability of NISQ devices to simulate quantum many-body dynamics through comparisons with exact solutions. The persistent and periodic oscillations in entanglement entropy, along with their analytical expressions, could potentially be used as a tool for calibrating quantum hardware and benchmarking quantum algorithms designed to compute this quantity.
The paper is organized as follows: In Sec. II, we define the model and the quench protocols considered in our simulations; we summarize some known properties of the quench dynamics. In Sec. III, we describe our simulation setups on the IBM-Q devices and present our simulation results, focusing on the Rényi entanglement entropy and twist order parameters. We conclude in Sec. IV with a summary and discussion. A derivation of the entanglement entropy is presented in Appendix A.
@@ II The Model and quench protocols
where $m$ indexes $N$ unit cells that are pairs of two sites, one on sublattice $A$ and one on $B$. The operator $a^{\dagger}_{m}$($b^{\dagger}_{m}$) creates a spinless fermion at the $A$($B$)-sublattice site in the $m$th unit cell, and $\mathcal{N}$ is the the number of fermions. The constants $2J$ and $2J^{\prime}$ are regarded as the intracell and intercell hopping amplitudes, respectively. The boundary term is affected by the parity of the number of fermions [30]; depending on evenness or oddness of $\mathcal{N}$, the boundary conditions of the fermionic chain are antiperiodic or periodic.
@@
which is equivalent to the ground state of the trivial flat-band Hamiltonian with $(J,J^{\prime})=(1,0)$. For a quench from $\left|{\psi_{\text{N}}}\right\rangle$, the post-quench state will reach a winding number $\nu=1$ at times $t_{\text{N}}=(n-\frac{1}{2})\frac{\pi}{4}$ with $n\in\mathbb{Z}^{+}$; for a quench from $\left|{\psi_{\text{S}}}\right\rangle$, the post-quench state will have a winding number $\nu=2$ at $t_{\text{S}}=(n-\frac{1}{2})\frac{\pi}{2}$ [21, 22]. Below we denote by $t_{\text{N}}^{*}$ and $t_{\text{S}}^{*}$ those time points at which a finite winding number occurs.
In real space, the periodic generation of the winding number manifests itself in the dynamics of intercell entanglement formation [37]. For the initial $\left|{\psi_{\text{N}}}\right\rangle$ state, the post-quench state consists of a set of nearest neighbor intercell entangled states at $t=t_{\text{N}}^{*}$; For the initial $\left|{\psi_{\text{S}}}\right\rangle$ state, the post-quench state develops entangled states connecting $A$ and $B$ sublattices in two next-nearest-neighbor cells (i.e. extended to three lattice spacings) at times $t=t_{\text{S}}^{*}$. The development of such local entanglement is evident from the time-dependent Wannier functions in the free-fermion system, as discussed in Appendix A.
@@ II.1 Entanglement entropy
We note that $S^{0\to 2}(t)=2S^{0\to 1}(t/2)$ [see Fig. 1(a)]. For a short closed chain of length $L=4$ with only two unit cells, the dynamical Rényi EE is reduced to $S^{0\to 1}(t)$. Technical details of the derivation of the dynamical Rényi EE are given in Appendix A.
We note that quasiparticle excitations, which typically contribute to a linear growth of the EE after a quench [39], are confined with a group velocity of zero in our flat-band model. Thus, the persistent oscillations in the EE indicate the absence of relaxation following quenches [40].
Figure 1: Exact data for quenches from a Néel state $\left|{\psi_{\text{N}}}\right\rangle$ (solid lines) and from a fully dimerized state $\left|{\psi_{\text{S}}}\right\rangle$ (dashed lines), driven by the flat-band Hamiltonian with $(J=0,J^{\prime}=1)$ and periodic boundary conditions. The panels from top to bottom show the evolution of (a) the Rényi entanglement entropy with a symmetric bipartition (given by Eqs. (12) and (13)), (b) the twist order parameter for $L=16$, and (c) the Berry phase for $L=16$, obtained by exact diagonalization.
@@ II.2 Twist operator
for a spin chain of length $L$, is a useful tool to characterize various topology properties. This unitary operator was first introduced in the Lieb-Schultz-Mattis (LSM) theorem [30, 41], which states that the amplitude, dubbed as the twist order parameter, given by
@@
in terms of the ground state of the XXZ spin chain can be seen as an indicator of a gapped or gapless excitation spectrum, depending on whether $z^{(q)}_{L}$ is nonvanishing or zero in the thermodynamic limit $L\to\infty$. A similar operator was introduced to spinless electron systems [42, 43]:
$U_{n}=\exp\biggl{(}i\frac{2\pi}{L}\sum_{j=1}^{L}jn_{j}\biggr{)}\,,$ (16)
where $n_{j}=S_{j}^{z}+1/2$ is the particle number operator at site $j$. Similar to the twist order parameter $z_{L}$ defined in Eq. (15), the ground-state expectation value
@@
is an order parameter that distinguishes an insulating phase and a gapless conducting phase at rational filling $n=p/q$ (where $p/q$ is an irreducible fraction) [44], depending on whether $z^{(q)}_{N}\neq 0$ or $z^{(q)}_{N}=0$ in the limit of $L\to\infty$ [45]. Both definitions of the twist operator are consistent with the periodic boundary conditions of the chain.
The application of the twist order parameters have been extended to the study of various gapped states. For instance, the amplitude $z_{L}$ was proposed as an order parameter to distinguish different types of valence-bond-solid(VBS)-like dimerization [31], even in random and inhomogeneous systems [46]. In the case of a spin-1/2 VBS state, a dimerized pattern of two-spin singlets (valence bonds) formed on odd links (i.e. intracell links) is a dimer phase called $(1,0)$-phase, with $z_{L}^{(1)}=1$ for $L\to\infty$. The fully dimerized initial state, $\left|{\psi_{\mathrm{S}}}\right\rangle$, we consider here corresponds to a $(1,0)$-phase. On the other hand, a dimer phase where two-spin singlets are formed on even links (i.e. intercell links), referred to as the $(0,1)$-phase, results in $z_{L}^{(1)}=-1$ for $L\to\infty$. This occurs because a $2\pi$ rotation induces a negative sign for a valence bond at the boundary $(j=L)$.
Here we apply the twist operator $U_{z}$ to the dynamical states in our quench protocols. In Fig. 1(b), we demonstrate results for time-dependent $z_{L}^{(1)}$ with $L=16$, obtained by exact diagonalization. We observe a periodic recurrence of the $(1,0)$-phase in the time-dependent $z_{L}^{(1)}$ during a quench from the fully dimerized state $\left|{\psi_{\mathrm{S}}}\right\rangle$. This dimerization is characterized by the peaks of $z_{L}^{(1)}$, with a peak value $[\cos(\pi/L)]^{L/2}$ for a finite chain of length $L$ with PBC [31]. Interestingly, the twist order parameter can also be applied to the quench from the Néel state $\left|{\psi_{\mathrm{N}}}\right\rangle$, where a dimerized phase emerges periodically at $t=t^{*}_{\text{N}}$. This dimerized phase is similar to a $(0,1)$-phase and can be identified by the negative peaks of $\text{Re}(z_{L})$, corresponding to $\text{Re}(z_{L})=-[\cos(\pi/L)]^{L/2}$. It is worth noting that, for the dynamical state $\left|{\psi_{\text{N}}(t)}\right\rangle$, the twist order parameter is not purely real, in contrast to $\left|{\psi_{\text{S}}(t)}\right\rangle$. This difference arises due to the breaking of inversion symmetry in $\left|{\psi_{\mathrm{N}}(t)}\right\rangle$.
Another application of the twist operator that we consider here is its connection to the Berry phase. In the free fermion model, the Berry phase represents the phase angle associated with the evolution of the quantum state as each particle undergoes a momentum shift of $2\pi/L$ induced by the twist operator. Moreover, the definition of the twist operator in Eq. (16) provides a means to calculate the expectation value of particle position, $X=\sum_{j}jn_{j}$, in an $N$-particle system of size $L$ with PBC at rational filling $N/L=p/q$ through [42, 45, 44]
$\langle{X}\rangle=\frac{L}{2\pi q}\text{Im}\ln[z_{N}^{(q)}]\,.$ (18)
Therefore, the Berry phase is directly related to the position expectation value [42, 45, 44]:
$\gamma_{L}=\text{Im}\ln[z^{(q)}_{N}]=\frac{2\pi q}{L}\langle{X}\rangle\,.$ (19)
This, in turn, establishes a link between the Berry phase and the macroscopic polarization of a many-electron system: $P=-e\gamma_{L}/2\pi$, where $e$ is the electron charge [42, 45].
We apply $U^{q}_{n}$ with $q=2$ to our dynamical states at half filling ($n=1/2$). As shown in Fig. 1(c) The Berry phase for $\left|{\psi_{\text{N}}(t)}\right\rangle$ jumps between $-\pi$ and $\pi$ at $t=t_{\text{N}}^{*}$, indicating a quantized dynamical Chern number $C=\pm 1$; this result agrees with previous studies based on the single-particle Hamiltonian in momentum space [24]. On the other hand, the Berry phase for $\left|{\psi_{\text{S}}(t)}\right\rangle$ remains zero for all $t$, reflecting that the dynamical state has inversion symmetry and the twist order parameter remains real.
@@ III.1 Time evolution
The default state on the IBM devices is a tensor product of $\left|{0}\right\rangle$ and the computational basis is also composed by the tensor product of the eigenstates of $\sigma^{z}$. We set up the initial states $\left|{\psi_{\text{N}}}\right\rangle$ and $\left|{\psi_{\text{S}}}\right\rangle$ for our quench protocol using the Pauli gate $\bm{X}$, the Hadamard gate $\bm{H}$ and the controlled-not (CNOT) gate $\bm{C_{X}}$, as shown in Fig. 2, where a two-spin singlet state is realized by a set of quantum gates $(\bm{C_{X}})(\bm{H}\otimes\bm{I})(\bm{X}\otimes\bm{X})$.
@@
For the post-quench Hamiltonian, the fully dimerized XX chain with $(J,J^{\prime})=(0,1)$, that we consider here, the quantum circuit for the time evolution operator $U(t)=\exp(-iHt)$ with an arbitrary time duration $t$ can be constructed without any Trotter error. The time evolution operator can be split into
@@
where the commutator identity $[\sigma_{j}^{x}\sigma_{j^{\prime}}^{x},\,\sigma_{j}^{y}\sigma_{j^{\prime}}^{y}]=0$ is applied. The quantum circuits required for the time evolution are then the gates to implement the operators $U^{xx}(t)=e^{-it\bm{X}\bm{X}}$ and $U^{yy}(t)=e^{-it\bm{Y}\bm{Y}}$, which can be transformed from the time evolution operator of an Ising spin pair $U^{zz}(t)=e^{-it\bm{Z}\bm{Z}}$. The circuit for the operator $U^{zz}(t)$ is composed of a rotation gate $\bm{R_{z}}(2t)$ and two $\bm{C_{X}}$ gates, as shown in Fig. 3 (a). We can convert the operator $U^{zz}(t)$ into $U^{xx}(t)$ using Hadamard gates via $\bm{H}\bm{Z}\bm{H}=\bm{X}$. Similarly, we implement $U^{yy}(t)$ by wrapping the time-evolution operator $U^{zz}(t)$ with Hadamard gates and $\bm{S}$-gates on each qubit, where $\bm{S}$ is a phase shift gate that rotates a $\pi/2$ radian about the $z$-axis, inducing $\bm{S^{\dagger}}\bm{X}\bm{S}=-\bm{Y}$. The circuits for $U^{xx}(t)$ and $U^{yy}(t)$ are shown in Fig. 3 (b) and Fig. 3 (c), respectively.
@@ III.2 The Rényi entanglement entropy
The key ingredient of the second-order Rényi entropy is the purity $\text{Tr}[\rho^{2}]$ of a given (reduced) density matrix. A recently developed measurement protocol employing statistical correlations of locally randomized measurements [47, 48, 49, 50] enables us to measure the purity with an adequate accuracy on current noisy quantum devices via classical postprocessing. This measurement procedure for our simulation of the Rényi entanglement entropy is as follows: (i) Generate a post-quench state, $\rho(t)$, of $L$ qubits at time $t$; (ii) Draw $N_{U}$ sets random unitaries of the form $U=u_{1}\otimes\cdots\otimes u_{L}$, where the $u_{i}$ are single-qubit Haar random unitaries sampled independently from the circular ensemble [49, 51]; (iii) Apply each set of $U$ to the state $\rho(t)$ and measure the the probabilities $P(\mathbf{s}_{I})$ of obtaining computational basis states $\left|{\mathbf{s}_{I}}\right\rangle=\left|{s_{1},\cdots,s_{N_{I}}}\right\rangle$ for the considered subsystem $I$ of $N_{I}$ qubits. The purity $\text{Tr}[\rho_{I}^{2}]$ for the reduced density matrix is then estimated using
@@
Figure 5: Results for the half-chain entanglement entropy for a chain with $L=4$ and PBC after quenches from the initial states $\left|{\psi_{\text{N}}}\right\rangle$ and $\left|{\psi_{\text{S}}}\right\rangle$, simulated on the IBM device ibm_mumbai. The dashed line is the analytical result given in Eq. (12).
@@
Figure 6: Results for the half-chain entanglement entropy for a chain with $L=8$ and PBC after quenches from the initial states $\left|{\psi_{\text{N}}}\right\rangle$ (a) and $\left|{\psi_{\text{S}}}\right\rangle$ (b), simulated on the IBM device ibm_montreal. The dashed lines are the analytical results given in Eqs. (12) and (13), respectively. The simulation results for $\left|{\psi_{\text{S}}}\right\rangle$ after applying error mitigation using the global depolarizing ansatz show an excellent match with the exact solution.
@@
For longer chains, there is however a serious decline in the accuracy of the simulation results, as shown in Fig. 6 for $L=8$. For the initial state $\left|{\psi_{\text{N}}}\right\rangle$, there is a entropy shift when compared to the exact solution. This offset, possibly due to some inherent error in the device, can be partially eliminated by aligning the entropy at $t=0$ with the exact value $S(t=0)=0$. We note that the shift mitigation approach has also applied in previous studies [53]. To eliminate the deviations beyond the constant shift, we apply a simple error mitigation strategy using the global depolarizing noise ansatz, along the lines of Ref. [15]. We assume that the entire quantum state of $L$ qubits under the noise is
@@
where $I$ denotes a subsystem of $N_{I}$ qubits. The error rate $p_{\text{tot}}$ can be extracted from $\text{Tr}[\rho^{2}]$ when the full density matrix for a pure state is considered since $\text{Tr}[\rho^{2}_{\text{exact}}]=1$. We have determined $p_{\text{tot}}$ for each time step to carry out the error mitigation for the time-dependent data. In addition, the mitigated data in Fig. 6 are adjusted to align the minimum to zero to correct negative entropies in the minima. The results for $\left|{\psi_{\text{S}}}\right\rangle$ after applying error mitigation clearly show an excellent match with the exact data, while the mitigated results for $\left|{\psi_{\text{N}}}\right\rangle$ with global depolarizing ansatz are merely passable. We note that the oscillation frequency of the dynamical EE for $\left|{\psi_{\text{N}}}\right\rangle$ is twice that of $\left|{\psi_{\text{S}}}\right\rangle$. As a result, there are more time points where the exact EE becomes zero (i.e., the reduced density matrix corresponds to a pure state) and more intervals where the EE remains close to zero. Randomized measurements are known to be less accurate for pure states [48, 49] because a pure (reduced) density matrix is not a probabilistic mixture of other states. This may explain why the dynamical EE associated with the initial state $\left|{\psi_{\text{N}}}\right\rangle$ shows a larger deviation from the exact solution compared to $\left|{\psi_{\text{S}}}\right\rangle$. The increased variance for $S\to 0$ is also evident in the raw data presented in Fig. 5.
Figure 8: A schematic plot of the heavy-hexagon lattice, widely adopted in the IBM quantum processors. Each circle represents a physical qubit. The blue circles form a ring layout and the red circles form a line layout that are selected in our simulations for a 12-qubit chain and a 8-qubit chain, respectively.
@@
The deviation from the exact solutions can have more than one source. For example, the accuracy of the estimated purity by randomized measurements depends crucially on the number of random unitaries, $N_{U}$, and the number of measurements per random unitary, $N_{M}$; as analyzed in [48, 49], the statistical errors scale with $1/\sqrt{N_{U}}$ and $1/N_{M}$, respectively. However, the primary source of error for our case may stem from the PBC imposed on the model, which requires connectivity between the first and last qubits via ancilla qubits and additional gates, thereby leading to higher error rates. For example, the circuit depth for the simulations of the chains with PBC is 40 (45) for the initial state $\left|{\psi_{\text{N}}}\right\rangle$ ($\left|{\psi_{\text{S}}}\right\rangle$), while the depth is reduced to 15 (19) when open chains with OBC are considered. We can indeed reproduce the results for $L=8$ with higher accuracy by considering the EE for a segment of bulk qubits (from the 3rd to the 6th qubit) in an open chain. The partition has two boundaries, resulting in the EE that is identical to that in the PBC case (cf. Appendix A). The results for the open chain displayed in Fig. 7 largely agree quantitatively with the exact data, even without applying error mitigation. After applying the global depolarizing error mitigation, the overall shape of the EE as a function of time, in particular around the entropy valleys, is corrected. However, we need to apply a constant shift to align the entropy at $t=0$ (the minimum entropy) to zero.